Write a quadratic equation having the given numbers as solutions.
8/5, 5/3
Possible answers:
A) 3x2 -40x + 25=0
B) 3x2 + 25x + 40=0
C) 3x2 + 49/5x + 8=0
D) 3x2 -49/5x + 8=0
E) 3x2 + 40x + 25=0

Answer by stanbon(75887)

Write a quadratic equation having the given numbers as solutions.
8/5, 5/3
Possible answers:
f(x) = k(5x-8)(3x-5)
f(x) = k(5x^2-49x+40)
Let k = (3/5)
Then
f(x) = 3x^2 - (3/5)49x + (3/5)40
f(x) = 3x^2 - 29.4x + 24
-----------------
Comment: The optional answers do not match the posted problem.
------
Cheers,
Stan H.
=====================
A) 3x2 -40x + 25=0
B) 3x2 + 25x + 40=0
C) 3x2 + 49/5x + 8=0
D) 3x2 -49/5x + 8=0
E) 3x2 + 40x + 25=0